I can not figure this out>>>> Thanks for your help
Marta leaves home at 9 am bicycling at a rate of 24 mi/h two hours later John leaves driving at a rate of 48 mi/h at what time will john catch up with Marta

Answer by Earlsdon(6103)

At 24 mph, Marta ought to consider competing in the Tour de France.
Using the formula: d = rt (distance = rate X time)
For Marta: d1 = (r1)(t1) = 24(t1)
For John: d2 = (r2)(t2) = 48(t2)
But Marta's time is 2 hours longer than John's time, so t1 = t2+2.
And when John catches up with Marta, they will both have traveled the same distance, so, d1 = d2
d2 = d1
48(t2) = 24(t2+2)
48t2 = 24t2 + 48
24t2 = 48
t2 = 2 and t1 = t2+2 = 4
Now Marta leaves at 9:00 am and takes 4 hours while John leaves at 11:00 am and takes 2 hours.
So they will meet up at 9:00 am + 4 hours = 1:00 pm or
11:00 am + 2 hours = 1:00 pm.